OpenStudy (anonymous):

evaluate integral 1/ (x(sqrt(4x^2+1))) dx

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ x \sqrt{4x^2 +1} } dx\]

myininaya (myininaya):

i would try a trig sub

OpenStudy (anonymous):

I went through it all, and i got an answer. We don't have an answer key so i checked on wolfram and it the answer seems off a bit. I'll write down what i did.

OpenStudy (anonymous):

beware of wolrams answers identical to yours via some tricky trickonometry

OpenStudy (anonymous):

first i let \[x = \frac{ 1 }{ 2 }\tan \theta \] \[dx =1/2 \sec^2\theta d \theta \] so when i simplified it i got \[\int\limits \sec \theta/ \tan \theta\] which also equals to \[\int\limits\limits \csc \theta\] i used table and got -ln|cscx +cotx| + C then i used trig to get theta into x my final answer is \[-\ln \left| \frac{ \sqrt{2x^{2}+1} }{ 2x } +\frac{ 1 }{ 2x }\right| + C\]

OpenStudy (anonymous):

is that correct?

myininaya (myininaya):

the only thing I don't like is sqrt(2x^2+1)

myininaya (myininaya):

should be sqrt(4x^2+1)

myininaya (myininaya):